OpenStudy (anonymous):

3log3 (6x-4) NB: log to base 3

OpenStudy (luffingsails):

Do you want help to simplify? What is the task?

OpenStudy (anonymous):

yes - i need help to simplify the expression

OpenStudy (anonymous):

3log3 (6x-4) =log3 (6x-4)^3

OpenStudy (anonymous):

why is that simpler?

OpenStudy (anonymous):

i am not saying it is wrong, just don't understand why one is any simpler than the other

OpenStudy (anonymous):

nothing more than that, unless converting into exponential form

OpenStudy (luffingsails):

I agree with Satellite. The original equation seems simplified to me.

OpenStudy (anonymous):

it depends ,somesime we need to synthesize no to analyze

OpenStudy (anonymous):

that depends on the problem given.

OpenStudy (anonymous):

in yours what he could do more?!

OpenStudy (anonymous):

hope you got my idea.

OpenStudy (luffingsails):

I get your idea but context of the task is important. Simplification has to be defined by the student asking. I think at least.

OpenStudy (anonymous):

3 log(3) * (6 x (-4)) = -34.3527303

OpenStudy (anonymous):

3 log(3) * (6 x (-4)) = -34.3527303 how do you get this number